# PatchAD: Patch-based MLP-Mixer for Time Series Anomaly Detection

### v3

@@ Abstract
Anomaly detection stands as a crucial aspect of time series analysis, aiming to identify abnormal events in time series samples. The central challenge of this task lies in effectively learning the representations of normal and abnormal patterns in a label-lacking scenario. Previous research mostly relied on reconstruction-based approaches, restricting the representational abilities of the models. In addition, most of the current deep learning-based methods are not lightweight enough, which prompts us to design a more efficient framework for anomaly detection. In this study, we introduce PatchAD, a novel multi-scale patch-based MLP-Mixer architecture that leverages contrastive learning for representational extraction and anomaly detection. Specifically, PatchAD is composed of four distinct MLP Mixers, exclusively utilizing the MLP architecture for high efficiency and lightweight architecture. Additionally, we also innovatively crafted a dual project constraint module to mitigate potential model degradation. Comprehensive experiments demonstrate that PatchAD achieves state-of-the-art results across multiple real-world multivariate time series datasets. Our code is publicly available.  [^1] https://github.com/EmorZz1G/PatchAD
@@ 1 Introduction
Time series anomaly detection is a pivotal aspect of data analysis, focused on identifying abnormal patterns within time-series data that significantly deviate from normal behavior Belay et al. (2023); Choi et al. (2021). This process becomes more complex in multivariate temporal anomaly detection, where it involves multiple channel dimensions. Driven by the rapid progression of large-scale sensing technologies and enhanced data storage capabilities, temporal anomaly detection has found widespread application in various sectors. For instance, in the field of the Internet of Things (IoT), it serves to monitor abnormal events in sensor data Yang et al. (2021). In healthcare sector, it plays a vital role in monitoring a person’s physiological data, thereby assisting in the early detection of health problems Cook et al. (2020).
However, extracting anomalies from extensive and complex temporal data presents significant challenges. First, accurately defining the representations of anomalies is difficult due to the varied and evolving nature of real-world anomalies. For instance, anomalies in machinery could manifest as either high or fluctuating temperatures under different scenarios. Second, labels for anomalous events in time-series data are probably absent or hard to acquire. Thirdly, the high dimensionality and temporal dependencies inherent in time-series data are intricate. While the typical approach involves classifying different time points as normal or anomalous, relying solely on a single time point is often not feasible.
Due to the powerful representational capacity of neural networks, numerous deep learning-based techniques Munir et al. (2019); Zhang et al. (2022) have been proposed for time series anomaly detection, demonstrating remarkable performance and attracting significant interest from the research community. While supervised and semi-supervised methods grapple with the scarcity of labeled data, unsupervised techniques Zhou et al. (2019); Su et al. (2019) have gained prominence due to their ability to operate without strict labeling requirements. While Xu et al. (2022) designed six different anomaly injection methods to train the model in a self-supervised approach, Carmona et al. (2021) utilized outlier exposure and mixup to construct a self-supervised task.
Reconstruction-based approaches, as typical and common unsupervised methods, involve learning representations of historical data and identifying samples that cannot be accurately reconstructed during testing as anomalies Park et al. (2018). This approach has experienced rapid development due to its clear concept and its proficiency in managing complex data representations. However, due to their robust reconstruction capabilities, these methods excel at recovering anomalous data, posing a challenge for reconstruction-based approaches. The core issue is the ineffectiveness of these methods in distinctly modeling normal and abnormal data, complicating their ability to accurately identify anomalies.
Recently, contrastive learning has been captivated by its straightforward, yet exceptional performance, proving effective in areas such as computer vision Chen et al. (2020). In time series anomaly detection, several methods have begun incorporating contrastive learning to enhance model representational capabilities. Nonetheless, these approaches often overlook the design of network structures tailored for contrastive learning. They either do not explicitly model inter-channel dependencies Yang et al. (2023) or rely solely on basic MLPs or Transformers that focus on temporal information Zhou et al. (2022), mainly considering point-level features Audibert et al. (2020). Nevertheless, poor training and inference speeds, architectural complexity, and a lack of lightweight features are some of the drawbacks of high-performing techniques. It was recently shown by Zeng et al. (2022) that MLP structures can also result in outstanding performance on time series tasks. Additionally, they neglected the model degradation issue in contrastive learning, meaning models might degrade or collapse during the training process.
In this paper, we propose PatchAD, a novel Patch-based MLP architecture designed for time series Anomaly Detection. PatchAD leverages contrastive learning to tackle the complexities of time series anomaly detection. Inspired by Yang et al. (2023), we hypothesize that normal time series points share similar latent patterns, indicating a strong correlation among normal time points. On the contrary, anomalies struggle to establish connections with other anomaly points or normal points. Hence, normal data exhibits consistent representations across different views, while anomalies struggle to maintain consistent representations. This underlying principle forms the core of our approach to anomaly detection.
@@
Specifically, we introduce a novel Multi-scale Patching and Embedding approach. This method decomposes the input data along the temporal dimension into non-overlapping patches, enabling the model to learn from features that contain richer semantic information, moving beyond simple point-level features. To address various time scales, different submodels are utilized, each focusing on different patch sizes. To capture diverse views of temporal data representation, we propose different Patch Mixer Encoders to learn the mixed relationships among inter-patches, intra-patches, and inter-channel connections. We utilize a shared MixRep Mixer to uniformly represent the feature space. To prevent the degradation, we not only construct diverse encoders but also introduce a Dual Project Constraint mechanism to prevent the model from settling on overly simplistic solutions, thus substantially boosting its representational capabilities.
The main contributions of our proposed PatchAD can be summarized as follows:
  • The novel PatchAD model, innovatively structured on contrastive learning framework, is adept at managing diverse anomaly types. This is achieved by creating Interpatch Encoders and Intra-patch Encoders, which generate both inter-patch and intra-patch perspectives.
  • Employing a strategy named Multi-scale Patching and Embedding, PatchAD integrates multi-scale subsequence modules constructed exclusively with MLPs. This design is intended to facilitate the learning of semantically richer features.
  • Dual Project Constraint mechanism is specifically designed to impose constraints on the model, reducing its susceptibility to simple, trivial solutions and thus increasing its expressive power.
  • PatchAD’s efficacy is underscored by comprehensive experimental results on eight datasets. These results highlight its superior performance compared to the SOTA time series anomaly detection methods.
@@
2 Preliminaries
In this section, we provide an initial exposition on the formal representation of time series anomaly detection, related methods, and contrastive learning. These are the foundational elements of the research problem and serve as the motivation for our work.
2.1 Problem Formulation
In this paper, we consider a multivariate time series of length $T$ in an unsupervised setting:
$\mathcal{X}=(x_{1},x_{2},\cdots,x_{T}),$ (1)
where $x_{t}\in\mathbb{R}^{C}$ represents a C-dimensional channel feature, such as data from C sensors or machines. The objective is to train a model using $\mathcal{X}_{train}$ and provide predictions $\mathcal{Y}_{test}=(y_{1},y_{2},\cdots,y_{T^{\prime}})$ on another sequence $\mathcal{X}_{test}$ of length $T^{\prime}$. Here, $y_{t}\in\{0,1\}$, where 1 signifies an anomaly at that time point and 0 indicates a normal point.
2.2 Time Series Anomaly Detection
@@
There exist various methods proposed for time series anomaly detection, encompassing statistical approaches, classical machine learning methods, and deep learning techniques. Statistical methods involve moving averages and the autoregressive integrated moving average (ARIMA) model. Classical machine learning methods encompass classification-based approaches like One-Class Support Vector Machine (OCSVM) Schölkopf et al. (2001) and Support Vector Data Description (SVDD) Tax and Duin (2004). Deep learning methods include DAGMM, OmniAnomaly Su et al. (2019), USAD Audibert et al. (2020), LSTM-VAE Park et al. (2018), and DeepAnT, which are trained on normal data and detect anomalies by inference using reconstruction errors. Some techniques Kim et al. (2023),Chen et al. (2022) have leveraged self-attention mechanisms and have demonstrated promising detection results. Recognizing the limitations imposed by the reconstruction-based assumption, research efforts like Li et al. (2021a) introduced sequence variational autoencoder model based on smoothness constraints, whereas Xu et al. (2021) proposed prior and sequence associations for temporal modeling, both achieving enhanced performance.
2.3 MLP Mixer for Time Series
@@
The MLP Mixer, initially proposed as a novel architecture for computer vision Tolstikhin et al. (2021), has gained significant attention in recent years. Recently, there has been analysis indicating that MLP Mixers can effectively handle sequential data and find applications in other domains, such as time series prediction. Specifically, works like Nie et al. (2023); Gong et al. (2023) have focused on prediction tasks, particularly for multi-step time series prediction, leveraging the capabilities of MLP Mixers to achieve improved forecasting results. Their experiments demonstrate that the MLP Mixer surpasses existing Transformer-based and LSTM-based methods in long sequence time-series forecasting (LSTF). To the best of our knowledge, there is currently an absence of MLP Mixer application in the time series anomaly detection.
2.4 Contrastive Learning
@@
Contrastive learning has garnered considerable attention across various domains for its capacity to train models by contrasting positive pairs against negative ones. The objective is to drive the model to bring positive pairs closer together in the learned representation space while pushing apart negative pairs. Additionally, methods like BYOL Grill et al. (2020) and SimSiam Chen and He (2021) have demonstrated advanced performance without explicitly constructing negative pairs. Recently, studies such as Eldele et al. (2021) and Pranavan et al. (2022) have explored the integration of contrastive learning into time series analysis. Yang et al. (2023), on the other hand, combines self-attention with contrastive learning, achieving state-of-the-art results in time series anomaly detection.
@@ 3.1 Overall Architecture
Figure 1 illustrates the overall structure of PatchAD. PatchAD adopts patch sizes at multiple scales. We exemplify its architecture using a single scale. PatchAD is an $L$-layer network, where each layer comprises distinct modules: MLP Mixers, Project head, and ReWeight, forming its structure. We have devised four distinct types of MLP Mixers: Channel Mixer, Inter Mixer, Intra Mixer, and MixRep Mixer, each dedicated to learning various channels, inter-patch, intra-patch relationships, and a unified representation space. The Project head serves to prevent the model from converging to trivial solutions, while the ReWeight module manages the weights across different layers.
The focal point of our design lies in the four distinct MLP Mixers. Inspired by Yang et al. (2023), we utilize various MLP Mixers to learn different representations from the input. Normal points share a common latent space across different views, while anomalies, being scarce and lacking specific patterns, struggle to share a coherent representation with normal points. Consequently, normal points exhibit minor discrepancies across different views, whereas anomalies showcase larger variations. PatchAD leverages inter-intra discrepancy to model this relationship.
@@ 3.2 Multi-scale Patching and Embedding
To enhance the temporal contextual representation in time series, we incorporated Transformer positional encoding. Acknowledging the significance of different sequence lengths in the time series analysis, we introduced Multi-scale Patching in PatchAD. This enables PatchAD to focus on features of varying sequence lengths. Moreover, incorporating multi-scale information helps compensate for information loss during the patching process. The input data with $C$ channels and length $T$ is transformed into patches of size $P$. It can be seen as dividing the time series data into $N$ non-overlapping blocks of patches, which is called $\operatorname{Patching}(\cdot)$. Thus, the original data of dimensions $C\times T$ is transformed into $C\times N\times P$.
@@
where $\operatorname{PE}(\cdot)$ is positional embedding.
PatchAD employs two value embeddings, i.e., $\operatorname{VE}(\cdot)$, facilitating the features from two distinct perspectives within the framework. And $\operatorname{VE}(\cdot)$ is a single linear layer.
@@ 3.3 Patch Mixer Encoder
PatchAD comprises $L$ layers of Patch Mixer layers. This encoder incorporates four distinct MLP Mixers to extract features across different dimensions. Figure 1(c) illustrates the MLP block we designed. It consists of two fully connected layers, a normalization layer, a GELU non-linear activation layer, and utilizes a residual connection from input to output. This can be represented as:
@@
    The Channel Mixer enables PatchAD to capture correlations among different channels.
@@
    The Inter Mixer empowers PatchAD to capture inter-patch representations, facilitating the learning of global expressions.
@@
    The Intra Mixer enables PatchAD to capture intra-patch representations, facilitating the learning of local expressions.
@@
    The MixRep Mixer allows PatchAD to embed two distinct views into the same representation space.
The diagram depicts PatchAD comprised of an inter-patch encoder and an intra-patch encoder, generating representations denoted as $\mathcal{N}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}\in\mathbb{R}^{C\times P\times D}$, representing the inter-patch and intra-patch views, respectively. It’s important to note that The Channel Mixer and The MixRep Mixer share weights, while The Inter Mixer and The Intra Mixer have their own weights. This design facilitates PatchAD in learning differences between views and prevents model degradation.
@@ 3.4 Dual Project Constraint
To prevent the model from converging to trivial solutions, each layer’s output requires to be constrained by Dual Project Head. Prior research suggests that employing a similar structure effectively prevents models from converging to trivial solutions. For instance, SimCLR Chen et al. (2020) uses a projection head that maps augmented views of an image into a latent space and promotes smoother convergence during training. The projection head comprises two layers of MLP connections, excluding non-linear and normalization layers. Hence, for both the inter-patch and intra-patch encoders, we can derive the projected representations $\mathcal{N}^{\prime}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}^{\prime}\in\mathbb{R}^{C\times P\times D}$. This computation can be represented as:

### v5

# PatchAD: A Lightweight Patch-Based MLP-Mixer for Time Series Anomaly Detection

@@ Abstract
Anomaly detection in time series analysis is a pivotal task, yet it poses the challenge of discerning normal and abnormal patterns in label-deficient scenarios. While prior studies have largely employed reconstruction-based approaches, which limits the models’ representational capacities. Moreover, existing deep learning-based methods are not sufficiently lightweight. Addressing these issues, we present PatchAD, our novel, highly efficient multiscale patch-based MLP-Mixer architecture that utilizes contrastive learning for representation extraction and anomaly detection. With its four distinct MLP Mixers and innovative dual project constraint module, PatchAD mitigates potential model degradation and offers a lightweight solution, requiring only 3.2MB. Its efficacy is demonstrated by state-of-the-art results across 9 datasets sourced from different application scenarios, outperforming over 30 comparative algorithms. PatchAD significantly improves the classical F1 score by 50.5%, the Aff-F1 score by 7.8%, and the AUC by 10.0%. The code is publicly available.  [^1] https://github.com/EmorZz1G/PatchAD
@@ 1 Introduction
Time series anomaly detection is a pivotal aspect of data analysis, focused on identifying abnormal patterns within time-series data that significantly deviate from expected behaviour [1, 2, 3]. With the rapid progression of large-scale sensing technologies and storage capabilities, temporal anomaly detection has found widespread application in various sectors. For instance, the Internet of Things (IoTs) serve to monitor abnormal events in sensor data [4, 5]. Unsupervised algorithms, capable of operating without labels, have gained wide popularity. These include reconstruction-based[6, 7], density-based[8, 9], contrastive learning[10, 11], autoregression-based[12, 13, 14], large language models (LLMs)[15] etc. However, extracting anomalies from extensive and complex temporal data presents significant challenges for different algorithms. Challenge 1: Accurately defining anomaly representations is challenging due to real-world anomalies’ diverse and dynamic nature. Effective anomaly detection requires models that can learn the interrelationships between various data channels. Challenge 2: There is still significant potential for improvement in terms of training and inference speed, model complexity, and the lack of lightweight features. Challenge 3: In an unsupervised setting, the challenge lies in establishing a reliable basis to model normal and anomalous data. Specifically, several algorithms have begun incorporating contrastive learning to enhance model representational capabilities. Nonetheless, these approaches often overlook the design of network structures and model degradation issue of contrastive learning. Additionally, they either do not explicitly model inter-channel dependencies [10] or rely solely on basic MLPs or Transformers that focus on temporal information [16], mainly considering point-level features [17].
On the other hand, we find that normal time series data points exhibit reduced information entropy, indicating higher similarity and correlation among them, whereas anomalies show increased entropy due to difficulties in establishing connections with other points. This difference aids our model in detecting anomalies, ensuring that normal data maintains consistent representations across different views, whereas anomalies do not. This fundamental principle forms the basis of our proposed anomaly detection method.
@@
In response to the above challenges and principles, we introduce PatchAD, a lightweight Patch-based MLP architecture for time series Anomaly Detection. Specifically, proposed Multi-scale Patching and Embedding decomposes the input data along the temporal dimension into non-overlapping patches, extracting semantically richer features beyond simple point-level characteristics. Besides, various submodels address various time scales, each focusing on different patch sizes. Patch Mixer Encoders are introduced to learn the complex relationships within and between patches and channels. A shared MixRep Mixer uniformly represents the feature space, while a Dual Project Constraint prevents the model from adopting overly simplistic solutions, thereby enhancing its representational capabilities. In detail, this work makes the following contributions:
Figure 1: Overall performances under different metrics.
Figure 2: Model sizes.
(1) State-of-the-art Anomaly Detection (Figure2): We propose a novel contrastive learning-based time series anomaly detection method. PatchAD achieves superior performance on 9 benchmark datasets, surpassing over 30 baselines across various evaluation metrics, with and without prior association.
(2) Lightweight and Efficient Design (Figure2): PatchAD employs a multi-scale patch-based MLP-Mixer architecture with different novel strategies, achieving remarkable efficiency and compactness compared to Transformer-based models.
(3) Comprehensive Analysis: An extensive ablation study validates the effectiveness of each component of PatchAD, supported by a theoretical analysis that elucidates the model’s underlying principles and design choices.
2 Preliminaries
2.1 Problem Formulation
In this paper, we consider a multivariate time series of length $T$ in an unsupervised setting: $\mathcal{X}=(x_{1},x_{2},\cdots,x_{T})$, $x_{t}\in\mathbb{R}^{C}$ represents a C-dimensional channel feature, such as data from C sensors or machines. The objective is to train a model using $\mathcal{X}_{train}$ and provide predictions $\mathcal{Y}_{test}=(y_{1},y_{2},\cdots,y_{T^{\prime}})$ on another sequence $\mathcal{X}_{test}$ of length $T^{\prime}$. Here, $y_{t}\in\{0,1\}$, where 1 signifies an anomaly at that time point and 0 indicates a normal point.
2.2 Time Series Anomaly Detection
@@
The time series anomaly detection methods encompass statistical approaches, classical machine learning methods, and deep learning techniques[18]. Statistical methods involve moving averages and the autoregressive integrated moving average (ARIMA) model. Classical machine learning methods encompass classification-based approaches like One-Class Support Vector Machine (OCSVM) [19] and Support Vector Data Description (SVDD) [20]. Deep learning methods include DAGMM, OmniAnomaly [7], USAD [17], LSTM-VAE [6], and DeepAnT [21], which are trained on normal data and detect anomalies by inference using reconstruction errors. Some techniques have leveraged self-attention mechanisms and have demonstrated promising detection results [22, 23]. While [24] designed six different anomaly injection methods to train the model in a self-supervised approach, [25] utilized outlier exposure and mixup to construct a self-supervised task. Recognizing the limitations imposed by the reconstruction-based assumption, research efforts like [26] introduced sequence variational autoencoder model based on smoothness constraints, whereas [27] proposed prior and sequence associations for temporal modeling, both achieving enhanced performance. Recently approaches based on diffusion models and Large language model have also been proposed [28, 15].
2.3 MLP Mixer & Contrastive Learning for Time Series Analysis
@@
The MLP Mixer initially proposed as a novel architecture for computer vision [29], has gained significant attention in recent years. Recently, there has been analysis indicating that MLP Mixers can effectively handle sequential data and find applications in other domains, such as time series prediction. Specifically, works like [30, 31, 32] have focused on prediction tasks, particularly for multi-step time series prediction, leveraging the capabilities of MLP Mixers to achieve improved forecasting results. Their experiments demonstrate that the MLP Mixer surpasses existing Transformer-based and LSTM-based methods in long sequence time-series forecasting (LSTF).
@@
Contrastive learning aims to drive the model to bring positive pairs closer together in the learned representation space while pushing apart negative pairs. Additionally, methods like BYOL [33] and SimSiam [34] have demonstrated advanced performance without explicitly constructing negative pairs. Recently, studies such as [11] and [35] have explored integrating contrastive learning into time series analysis. [10], on the other hand, combines self-attention with contrastive learning, achieving state-of-the-art results in time series anomaly detection.
@@ 3.1 Overall Architecture
Figure 3 illustrates the overall structure of PatchAD, which adopts patch sizes at multiple scales. We exemplify its architecture using a single scale. PatchAD is an $L$-layer network, where each layer comprises distinct modules: MLP Mixers, Project head, and ReWeight, forming its structure. We have devised four distinct types of MLP Mixers: Channel Mixer, Inter Mixer, Intra Mixer, and MixRep Mixer, each dedicated to learning various channels, inter-patch, intra-patch relationships, and a unified representation space. The Project head serves to prevent the model from converging to trivial solutions, while the ReWeight module manages the weights across different layers.
The focal point of our design lies in the four distinct MLP Mixers. We utilize various MLP Mixers to learn different representations from the input. Normal points share a common latent space across different views, while anomalies, being scarce and lacking specific patterns, struggle to share a coherent representation with normal points. Consequently, normal points exhibit minor discrepancies across different views, whereas anomalies showcase larger variations. PatchAD leverages inter-intra discrepancy to model this relationship.
Our anomaly detection approach, PatchAD, leverages the principle that normal time series points, characterized by low information entropy, exhibit consistent representations across diverse perspectives. In contrast, anomalies with higher entropy struggle to maintain such consistency (Figure 3(c)). PatchAD employs an Inter-Patch Encoder to capture detailed features and an Intra-Patch Encoder to learn generalized trends. This combined approach enables a comprehensive understanding of normal and anomalous representations, achieving effective anomaly detection.
@@ 3.2 Multi-Scale Patching and Embedding
To enhance the temporal contextual representation in time series, we incorporated Transformer positional encoding. Acknowledging the significance of different sequence lengths in the time series analysis, we introduced Multi-scale Patching in PatchAD. This enables PatchAD to focus on features of varying sequence lengths. Moreover, incorporating multi-scale information helps compensate for information loss during the patching process. The input data with $C$ channels and length $T$ is transformed into patches of size $P$. It can be seen as dividing the time series data into $N$ non-overlapping patches blocks, called $\operatorname{Patching}(\cdot)$. Thus, the original data of dimensions $C\times T$ is transformed into $C\times N\times P$.
@@
where $\operatorname{PE}(\cdot)$ is positional embedding. PatchAD employs two value embeddings, i.e., $\operatorname{VE}(\cdot)$, facilitating the features from two distinct perspectives within the framework. And $\operatorname{VE}(\cdot)$ is a single linear layer.
@@ 3.3 Patch Mixer Encoder
PatchAD comprises $L$ layers of Patch Mixer layers. This encoder incorporates four distinct MLP Mixers to extract features across different dimensions. The MLP Mixer consists of two fully connected layers, a normalization layer and a GELU non-linear activation layer, and utilizes a residual connection from input to output. This can be represented as:
@@
    Channel Mixer captures correlations among different channels.
@@
    Inter Mixer captures inter-patch representations, improving detailed expressions.
@@
    Intra Mixer captures intra-patch representations, improving generalized expressions.
@@
    MixRep Mixer embeds two distinct views into the same representation space.
The diagram depicts PatchAD comprised of an inter-patch encoder and an intra-patch encoder, generating representations denoted as $\mathcal{N}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}\in\mathbb{R}^{C\times P\times D}$, representing the inter-patch and intra-patch views, respectively. It’s important to note that The Channel Mixer and The MixRep Mixer share weights, while The Inter Mixer and The Intra Mixer have their own weights. This design facilitates PatchAD in learning differences between views and prevents model degradation. In the end, two separate simple MLPs are utilized to reconstruct the original data, which are then combined.
$\hat{\mathcal{X}}_{1}=\operatorname{Rec1}(\mathcal{N}),\quad\hat{\mathcal{X}}_{2}=\operatorname{Rec2}(\mathcal{P}),\quad\hat{\mathcal{X}}=\hat{\mathcal{X}}_{1}+\hat{\mathcal{X}}_{2}.$ (4)
@@ 3.4 Dual Project Constraint
To prevent the model from converging to trivial solutions, each layer’s output requires to be constrained by Dual Project Head. Prior research suggests that employing a similar structure effectively prevents models from converging to trivial solutions. For instance, SimCLR [36] uses a projection head that maps augmented views of an image into a latent space and promotes smoother convergence during training. The projection head comprises two layers of MLP connections, excluding non-linear and normalization layers. Hence, for both the inter-patch and intra-patch encoders, we can derive the projected representations $\mathcal{N}^{\prime}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}^{\prime}\in\mathbb{R}^{C\times P\times D}$. This computation can be represented as: